OpenStudy (anonymous):

Solve each quadratic equation. Show your work. 4x2 = 100

Mehek (mehek14):

4 times 2 = 100? is that the question?

OpenStudy (welshfella):

4 x^2 = 100

OpenStudy (anonymous):

yeah

OpenStudy (texaschic101):

4x^2 = 100 -- divide both sides by 4 x^2 = 100/4 x^2 = 25 --- now take the sq rt of both sides...this cancels out the ^2 on the right side. x = sq rt 25 x = 5 check.. 4(5^2) = 100 4(25) = 100 100 = 100 (correct) so x = 5

OpenStudy (anonymous):

oh ok thakx

OpenStudy (texaschic101):

sure thing :)